OpenStudy (anonymous):

ANSWERS GET MEDALS!! Find the Domain 1. f(z)=√((1-z)/(z+2)) 2. √-x^2-15x-54 3. y=√-x^2+12x-27

OpenStudy (anonymous):

1) First: you can't divide by 0, so z+2 can not be 0, then z not equal to (-2) second condition - expression under sq. rt >=0

OpenStudy (anonymous):

for that - put 2 systems together: a) 1-z>=0 and z+2>=0 b) 1-z<=0 and z+2<=0 solve it

OpenStudy (anonymous):

should get: z : (-2,1]

OpenStudy (anonymous):

ok thank you thats what i got for that one as well what about the other two?

OpenStudy (anonymous):

2) \[=\sqrt{-(x+9)(x+6)}\] now you have to solve it that the expression under (all!) will be >=0 which means that one of the multiples needs to be negative

OpenStudy (anonymous):

or: a) x+9<=0 and x+6 >=0 b) x+9 >=0 and x+6<=0 get your answer & let's compare

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

9,6

OpenStudy (anonymous):

i got: x: [-9,-6] and another set is empty check signs in your solution :)

OpenStudy (anonymous):

3)\[\sqrt{-(x-9)(x-3)}\] need to do the same as in 2) do you want to try & will compare notes?

OpenStudy (anonymous):

[3,9]..?

OpenStudy (anonymous):

should have: a) x-9>=0 and x-3<=0 b) x-9 <=0 and x-3 <=0 You are right!!

OpenStudy (anonymous):

sweet!! thanks man!!

OpenStudy (anonymous):

:))